EINSTEIN A COEFFICIENTS FOR VIBRATION-ROTATIONAL TRANSITIONS OF NO

\underline{M. GUTIÉRREZ}, School of Chemistry and Biochemistry, Georgia Institute of Technology, Atlanta, GA 30332, USA; J. F. OGILVIE, Escuela de Química and CELEQ, Universidad de Costa Rica, San José 2060, Costa Rica.

Using an algebraic approach^{a} with software for symbolic computation, we calculated the Einstein A coefficients for vibration-rotational transitions with Δv = 1 and 2 for NO in its electronic ground state, ^{2}Π, in substates both Ω = 1/2 and 3/2, up to v = 10. These values will be applicable in an analysis of the chemiluminescence of NO resulting from exothermic chemical reactions in the gaseous phase.

aJ. F. Ogilvie, The Vibrational and Rotational Spectrometry of Diatomic Molecules, Academic Press, London UK, 1998